# Xi+: Uncertainty Supervision for Robust Speaker Embedding

### v2

@@ 2.1 Recap of Xi-vector embedding with Uncertainty
$\boldsymbol{\phi}$ $=\frac{\sum_{t=1}^{T}\mathbf{L}_{t}\mathbf{z}_{t}+\mathbf{L}_{p}\mathbf{z}_{p}}{\mathbf{L}_{s}},$ (3)
@@ 2.2.1 Temporal Modeling
The original xi-vector architecture predicts the precision matrix using fully connected layers, which only capture relationships along the channel dimension. We hypothesize that temporal dependencies are also informative for uncertainty modeling. To this end, we introduce a Transformer encoder containing a single layer with NN attention heads within the pooling layer  [^2] NN is set to 8 in this experiment.. This newly added module serves two primary purposes: enhancing the non-linear modeling capacity of the network, and enabling the incorporation of temporal contextual information across frames.
@@ 3.1 Training details
where $\kappa$ is used to control loss weight and is dynamically adjusted based on the current epoch epep to ensure that $\mathcal{L}_{\text{SVL}}$ gradually contributes as training progresses:

### v3

@@ 2.1 Recap of Xi-vector embedding with Uncertainty
$\boldsymbol{\phi}$ $=\frac{\sum_{t=1}^{T}\mathbf{L}_{t}\mathbf{z}_{t}+\mathbf{L}_{p}\mathbf{z}_{p}}{\mathbf{L}},$ (3)
@@ 2.2.1 Temporal Modeling
The original xi-vector architecture predicts the precision matrix using fully connected layers, which only capture relationships along the channel dimension. We hypothesize that temporal dependencies are also informative for uncertainty modeling. To this end, we introduce a Transformer encoder containing a single layer with $N$ attention heads within the pooling layer  [^2] NN is set to 8 in this experiment.. This newly added module serves two primary purposes: enhancing the non-linear modeling capacity of the network, and enabling the incorporation of temporal contextual information across frames.
@@ 3.1 Training details
where $\kappa$ is used to control loss weight and is dynamically adjusted based on the current epoch $ep$ to ensure that $\mathcal{L}_{\text{SVL}}$ gradually contributes as training progresses: